# Exact holomorphic solution of the Hodge-dual minimal surface: a confining quantum QCD string

### v1

@@ 7 Reparametrization and additivity
Compatibility with the MM kernel. The right–hand side of the MM equation [5, 6] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
@@
so the kernel is parametrization–invariant. Therefore the canonical standardization to $0,2\pi]$ used above is legitimate and (7.4) guarantees that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\uplus\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [7, 2]).
@@ 8 Confinement theorem
Let $G$ be the Douglas–Gram matrix of $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. as we have proven in the previous section, the Hodge–dual area and the Euclidean minimal area satisfy
@@
Both functionals, Douglas’s $\sum\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, have yet to be minimized over the reparametrization of the bounding loop $C(\theta)\Rightarrow C(f(\theta)),f^{\prime}>0$. As the minimization of just the two highest eigenvalues, compared to that of the trace, is attained at different parametrizations, we can only state that $\sigma_{eff}=2\sqrt{2}\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}$. However, the lower inequality still holds $\sigma_{eff}\geq\sqrt{2}$, as for every parametrization $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$.Therefore, the minimum in the denominator is less than or equal to twice that in the numerator, proving the lower bound quoted in confinement theorem.
9 Quantization
Our area functional is equivalent to the string theory with action $S_{0}[f_{i}]=\int_{\mathbb{D}}\!\mathrm{d}^{2}z\,(\nabla f_{i})^{2}+\int_{\mathbb{D}}\!\mathrm{d}^{2}z\,u_{i}\,\partial_{\bar{z}}f_{i}$, where $f_{1,2}(z)$ are the basis functions introduced in Section 4. On the boundary, we have linear boundary conditions of the form: $S_{\partial}[f_{i},\rho]=\oint\!\mathrm{d}\theta\,\rho_{i}(\theta)[\Re f_{i}\left(e^{\mathrm{i}\theta}\right)-Q_{i}(C(\theta))]$ with $Q_{1}(C)=(C_{1}+C_{2})/\sqrt{2},Q_{2}(C)=(C_{3}+C_{4})/\sqrt{2}$. Integrating out $f_{i}$ reproduces a Douglas-type boundary action $\mathcal{D}[QC]=\operatorname{tr}PG$. The functional integral over $f_{i}$ defines the quantum Hodge–dual string. The area derivative’s self-duality stems from local Hodge duality and holds quantum-mechanically. The quadratic action ensures quantum Schwinger–Dyson equations match classical E-L relations. Thus, the expectation value $\left\langle\delta S/\delta\sigma\right\rangle$ is self-dual, and the quantum string obeys the classical loop equations. Integrating out $f_{i}$ gives a Gaussian integral $Z[C;P]\!\propto\![\det{}^{\prime}(K_{P})]^{-1/2}\exp\left(-\,2\sqrt{2}\,\kappa\operatorname{tr}PG\right)$, where $K_{P}=P\,H\,\partial_{\theta}P$ has a universal spectrum. Hence $\det{}^{\prime}(K_{P})$ is a constant, and the effective action is classical: $\Gamma[C;P]=2\sqrt{2}\,\operatorname{tr}PG+\text{const}$. Since $S_{\chi}[C;P]=2\sqrt{2}\,\operatorname{tr}(PG)$ is spectral, $SO(3)_{+}$ rotations of $P$ are a gauge orbit. The integration factors out, and $Z[C]$ is evaluated on the spectral projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$: $Z[C]\;\propto\;\exp\left(-\,2\sqrt{2}\,\kappa(\lambda_{3}+\lambda_{4})\right)$. The gauge orbit carries no dynamics; $P_{\ast}$ captures the full invariant content.
10 Application to QCD
The known UV corrections (perimeter/cusp terms) to the QCD Wilson loop are sub-leading at large loop size $L\to\infty$. As we argue in detail in [2], due to the Leibniz property of the loop diffusion operator $\mathcal{L}$, the transformation $W[C]\Rightarrow W[C]\exp\left(-\kappa S[C]\right)$ leaves the MM loop equation[7] $\mathcal{L}W[C_{1}\uplus C_{2}]=W[C_{1}]\otimes W[C_{2}]$ invariant if $\mathcal{L}S[C]=0$ and $S[C]$ are additive on self-intersecting loops $S[C_{1}\uplus C_{2}]=S[C_{1}]+S[C_{2}]$. Our SD/ASD area satisfies both of these conditions (Euclidean area satisfies only additivity). SD or ASD areas satisfy the equation $\mathcal{L}S=0$ by virtue of the loop space Bianchi identity [2], just as the SD Yang–Mills field satisfies the equations of motion. Given the perturbative solution $W_{0}[C]$ we have the general solution as $W[C]=W_{0}[C]\exp\left(-\kappa S[C]\right)$, where the string mass scale $\kappa$ must be fixed as a universal number times the QCD mass scale by matching the perturbative region of small loops.
Crucially, the MM equation requires factorization of the loop derivative $\mathcal{L}W$, not the functional itself. As shown in [7], while the perturbative $W_{0}$ retains non-trivial correlations between sub-loops, the graphs connecting these sub-loops cancel upon application of $\mathcal{L}$, leaving exactly the factorized contact term on the right hand side. Our ansatz remains consistent because the geometric action is strictly additive, $S[C_{1}\oplus C_{2}]=S[C_{1}]+S[C_{2}]$. Combined with the zero-mode property $\mathcal{L}S=0$, the factor $\exp\left(-\kappa S\right)$ cancels identically from both sides of the equation, allowing the confining factor to dress $W_{0}$ without constraining its internal correlations.
@@ 11 Conclusion
This Hodge–dual minimal area $S[C]$ enters the non-perturbative factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop, solving the MM equation in QCD [5, 6, 7].
The primary physical implication is a first-principles geometric proof of confinement, as stated in our Confinement Theorem (Theorem 8.1). This theorem provides an explicit, shape-dependent formula for the effective string tension $\sigma_{\mathrm{eff}}[C]$ (defined as a ratio of $S[C]$ to the Euclidean area $A[C]$). This parameter depends on the loop shape and is bounded from below as $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$. This solution lays the groundwork for a quantitative theory of the high-level hadron spectrum.

### v2

@@ 8 Reparametrization and additivity
Compatibility with the MM kernel. The right–hand side of the MM equation [MMEq79, MM1981NPB] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
@@
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and (8.4) guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\uplus\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [Mig83, migdal2025SQYMflow]).
Let us stress that this factorization takes place independently of the splitting points $\theta_{1},\theta_{2}$. Thus, for every $\theta_{1},\theta_{2}$, the product of two factors $\exp\left(-\kappa S[C_{1}\right)$ and $\exp\left(-\kappa S[C_{2}]\right)$ exactly equals the factor $\exp\left(-\kappa S[C]\right)$ on the left side of the MM equation; thus, it can be taken out of the integral, after which it cancels on both sides of the MM equation.
@@ 9 Confinement theorem
Let $G$ be the Douglas–Gram matrix of $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. as we have proven in the previous section, the Hodge dual area and the Euclidean minimal area satisfy
@@
Both functionals, Douglas’s $\sum\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, have yet to be minimized over the reparametrization of the bounding loop $C(\theta)\Rightarrow C(f(\theta)),f^{\prime}>0$. As the minimization of just the two highest eigenvalues, compared to that of the trace, is attained at different parameterizations, we can only state that $\sigma_{eff}=2\sqrt{2}\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}$. However, the lower inequality still holds $\sigma_{eff}\geq\sqrt{2}$, as for every parametrization $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$.Therefore, the minimum in the denominator is less than or equal to twice that in the numerator, proving the lower bound quoted in confinement theorem.
10 Master Field and the Meaning of the Minimal Surface in Planar QCD
Planar QCD represents the large-$N_{c}$ limit of a Quantum Field Theory defined by gauge fields, summing over all planar graphs to describe fluctuations around a (yet unknown) vacuum state. However, when formulated in terms of the Wilson loop observable $\Psi[C]=\langle\frac{1}{N}\mathrm{Tr}\mathcal{P}\exp\left(\oint_{C}Adx\right)\rangle$, the theory undergoes a fundamental shift: it ceases to be a quantum theory and becomes a classical field theory in loop space.
In the large $N_{c}$ limit, due to factorization $\langle W(C_{1})W(C_{2})\rangle=\langle W(C_{1})\rangle\langle W(C_{2})\rangle+O(1/N^{2})$, the functional $\Psi[C]$ becomes the Master Field, satisfying a deterministic nonlinear equation— the MM loop equation.
It is crucial to understand that in this framework, we are not “quantizing a string theory” in the traditional sense. Rather, we are solving a classical field equation in loop space. The introduction of the Hodge-dual minimal surface in the auxiliary space $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ is a method to construct this solution.
This approach mirrors the logic of the Eguchi-Kawai (EK) reduction [Das:1987dg]. In the EK model, quenched “momenta” or twisted boundary conditions are not introduced as inherent physical degrees of freedom of the vacuum, but as mathematical instruments designed to satisfy the loop equations of the planar theory on a single lattice site. Similarly, in our construction, the only physical requirement is the satisfaction of the MM equation; the string analogy applies strictly insofar as it serves this goal.
There is no claim of a direct physical connection between the microscopic QCD gluon variables and the embedding coordinates $X_{\mu}^{A}$ in the 16-dimensional auxiliary space. Unlike the AdS/CFT correspondence, we do not postulate a full quantum equivalence between planar QCD and a fundamental string theory in a higher-dimensional bulk. Instead, we construct the specific additive zero mode of the loop equation, in the same spirit as one solves for the spontaneous magnetization in a ferromagnet: identifying the non-perturbative classical configuration that creates the confining background.
Still, the technology of String Theory, if not its complete physical contents, can be quite useful in elaborating this solution of planar QCD.
11 Topological String with Nonlinear Endpoint Dynamics
Our area functional is equivalent to the extremal value of the string theory action
$S_{0}[f_{i}]$ $=\int_{\mathbb{D}}d^{2}z\,(\nabla f_{i})^{2}+\int_{\mathbb{D}}d^{2}z\,u_{i}\,\partial_{\bar{z}}f_{i},$ (11.1)
where $f_{1,2}(z)$ are the basis functions introduced in Section 4. On the boundary, we have linear boundary conditions of the form:
$S_{\partial}[f_{i},\rho]$ $=\oint d\theta\,\rho_{i}(\theta)\left[\Re f_{i}\left(e^{i\theta}\right)-Q_{i}(C(\theta))\right],$ (11.2)
with $Q_{1}(C)=(C_{1}+C_{2})/\sqrt{2}$ and $Q_{2}(C)=(C_{3}+\chi C_{4})/\sqrt{2}$, where $\chi=\pm 1$ is the Hodge parity. Integrating out the auxiliary fields $f_{i}$ reproduces the Douglas-type boundary action $\mathcal{D}[QC]=\mathrm{Tr}(PG)$.
This system describes a topological field theory: the action reduces entirely to boundary values as a consequence of the imposed Cauchy-Riemann equations. Physically, this action represents the net area of two conformal maps, $f_{1}(\mathbb{D})$ and $f_{2}(\mathbb{D})$, with no remaining bulk degrees of freedom. In string theory language, there are no vibrations of the string body; there are only the nonlinear dynamics of its endpoints determined by the loop shape. In this sense, it is analogous to 2D quantum gravity, where the Einstein-Hilbert action reduces to the Euler characteristic, independent of the local geometry in the bulk.
The self-duality of the area derivative stems from local Hodge duality and holds quantum-mechanically. The quadratic nature of the action ensures that the quantum Schwinger–Dyson equations match the classical Euler-Lagrange relations. Thus, the expectation value $\langle\delta S/\delta\sigma\rangle$ is self-dual, and the quantum string obeys the classical loop equations.
Integrating out $f_{i}$ yields a Gaussian integral:
$Z[C;P]$ $\propto[\det{}^{\prime}(K_{P})]^{-1/2}$
$\times\exp\left(-2\sqrt{2}\,\kappa\mathrm{Tr}(PG)\right),$ (11.3)
where $K_{P}=P\,H\,\partial_{\theta}P$ has a universal spectrum. Hence $\det{}^{\prime}(K_{P})$ is a constant, and the effective action is classical: $\Gamma[C;P]=2\sqrt{2}\,\mathrm{Tr}(PG)+\text{const}$.
According to our min-max-min principle, which is required to preserve the factorization of the product of dressing factors $\exp\left(-\kappa S[C]\right)$ on the right side of the MM equation, we do not integrate over the “moduli space” (the rotation matrix $R$). Instead, we select the optimal rotation that brings the exponent to its maximal value:
$Z[C]$ $\propto\exp\left(-2\sqrt{2}\,\kappa\max_{R\in SO(4)}\mathrm{Tr}(R^{T}PRG)\right)$
$=\exp\left(-2\sqrt{2}\,\kappa(\lambda_{3}+\lambda_{4})\right).$ (11.4)
This corresponds to setting an optimal isometry of the boundary condition rather than averaging over these isometries. In terms of the partition function $Z=\exp\left(-\kappa S[RC]\right)$, our prescription selects the minimal value, rather than performing a saddle point integration around the maximal value of this factor.
Let us stress that we are solving the classical loop equations of large-$N_{c}$ QCD rather than quantizing a string theory. The internal degrees of freedom found in non-critical string theory (such as Polyakov’s Liouville mode) do not apply here, as our theory is based on the classical dynamics of Wilson loop trace functionals in planar QCD. We utilize string theory as a mathematical tool, rather than as a physical object describing the condensation of electric flux in the QCD vacuum.
In our picture, the fluctuations of this electric flux are described by planar graphs for gluons in another factor $W_{pert}[C]$. The confining factor $Z[C]$ only reflects the mean properties of the vacuum, frozen in the large $N_{c}$ limit. The optimal rotation $R_{*}$ described in section 5.1, aligning the SD/ASD projector with the highest plane of the Gram basis determines the boundary conditions for the holomorphic string:
$\Re f_{i}\left(e^{i\theta}\right)=Q_{i}(R_{*}C(\theta)).$ (11.5)
For the ASD string, the reflected coordinate $C_{4}$ is used in $Q_{i}(RC)$.
12 String Scale $\kappa$ and Gluon Condensate
The string mass scale $\kappa$ is not fixed by the MM equation itself. As we have seen, the confining factor $\exp\left(-\kappa S[C]\right)$ passes through the equation without correlation with the perturbative solution $W_{pert}[C]$. This factor reflects the non-perturbative properties of the QCD vacuum rather than the perturbative fluctuations of the gluon field.
The fundamental non-perturbative property of the QCD vacuum is the gluon condensate [Shifman:1978bx]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
At small distances $|x_{12}|=2r$, the Operator Product Expansion (OPE) in asymptotically free QCD is dominated by the leading conformal term in $W_{pert}[C]$:
$\frac{\delta^{2}W_{pert}[C]}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}$ $\propto\langle\mathrm{Tr}F_{\mu\nu}(x_{1})U_{0,\pi}F_{\lambda\rho}(x_{2})U_{\pi,2\pi}\rangle$
$\to\frac{\alpha_{eff}}{4\pi}\frac{I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}}{|x_{12}|^{4}}W_{pert}[C],$ (12.1)
where $I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda}$ with $\hat{x}=x_{12}/|x_{12}|$. We project this perturbative singularity out by contracting with the tensor $\Pi_{\mu\nu\lambda\rho}=\tfrac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})$. Simple tensor algebra confirms that the leading term vanishes: $\mathrm{Tr}(\Pi I\otimes I)=0$.
The remaining term is the gluon condensate contribution. This term involves the renormalization factor $Z^{2}$ associated with the two cusps created at the points of differentiation:
$\Pi_{\mu\nu\lambda\rho}\mathrm{Tr}F_{\mu\nu}(1)U_{1,2}F_{\lambda\rho}(2)U_{2,1}\to$
$-\langle(G^{a}_{\mu\nu})^{2}\rangle W_{pert}[0]Z^{2}.$ (12.2)
This constant contribution must be balanced by the second area derivative of our Hodge-dual factor in the ansatz $W=W_{pert}e^{-\kappa S}$. Using the fact that the first derivative $\delta S/\delta\sigma$ vanishes for the circle due to symmetry, we find:
$W_{pert}[C]\frac{\delta^{2}\exp\left(-\kappa S[C]\right)}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}\to$
$-\kappa W_{pert}[0]\Pi_{\mu\nu\lambda\rho}\frac{\delta^{2}S[C]}{\delta\sigma_{\mu\nu}(0)\delta\sigma_{\lambda\rho}(\pi)}.$ (12.3)
Crucially, there is no $Z$ factor in this geometric term; it depends only on the renormalization of the perturbative loop $W_{pert}[0]$ which appears as a common prefactor.
The second area derivative is computed as the second variation of $S=2\sqrt{2}(\lambda_{3}+\lambda_{4})$:
$\frac{\delta^{2}S[C]}{\delta\sigma^{2}}\propto\mathrm{Tr}(\Pi B\otimes B)\sum_{i<j}\frac{\left(\sum_{n>1}\frac{1}{n}\right)^{2}}{\lambda_{j}-\lambda_{i}}.$ (12.4)
The harmonic sum $\sum_{n>1}1/n\to\log n_{max}$ arises from the singular Hilbert kernels in the area derivatives. The result is:
$\frac{\delta^{2}S[C]}{\delta\sigma^{2}}=\frac{a(\log n_{max})^{2}}{r^{2}}.$ (12.5)
The sign matches the gluon condensate term, while the magnitude depends on the radius.
Remark 12.1.
The divergence scaling as $1/r^{2}$ arises from the second area derivative as a direct consequence of the scaling property of our Hodge minimal area. This is a nonlinear, nonlocal, homogeneous functional of degree two. Every area derivative involves two functional derivatives with respect to $\dot{C}$, lowering the degree by two. Thus, the second area derivative has degree minus two, inevitably scaling as $1/r^{2}$. The logarithmic divergence comes from the high-frequency spectral sum.
The proper physical procedure is to match these terms at the crossover scale. In the matching region, we equate the renormalized OPE result with the geometric derivative:
$Z^{2}\langle(G^{a}_{\mu\nu})^{2}\rangle=\frac{\kappa a(\log n_{max})^{2}}{r^{2}}\exp\left(-2\sqrt{2}\pi r^{2}\kappa\right).$ (12.6)
where the constant $a\sim 1$ is calculable from our solution, and renormalization constant $Z$ for the area derivative operator applied to $W_{pert}[C]$ is calculable by the methods of [KorchemskyRadyushkin1987]. Presumably, the renormalization factor $Z^{2}$ on the LHS compensates the UV divergence $(\log n_{max})^{2}$ on the RHS, rendering the relation finite and renormalization-group invariant.
The function $f(z)=ze^{-z}$ (where $z=2\sqrt{2}\pi r^{2}\kappa$) reaches its maximum at $z=1$. This implies that no solutions exist outside this parameter region, suggesting a critical point. We identify this maximum as the natural matching point:
$r_{c}^{2}=\frac{1}{2\sqrt{2}\pi\kappa};$ (12.7)
$\kappa^{2}=\frac{e}{2\sqrt{2}a\pi}\left(\frac{Z}{\log n_{max}}\right)^{2}\left\langle(G^{a}_{\mu\nu})^{2}\right\rangle.$ (12.8)
The factors of $N_{c}$ cancel, consistent with the planar limit. This matching fixes the string tension $\kappa$ in terms of the gluon condensate $\langle G^{2}\rangle$.
Our association of $\kappa$ with the gluon condensate supports the physical picture of condensed electric flux originally conjectured by Nambu [Nambu:1974zg]. One may view the term $\kappa S[C]$ as the action cost of a field configuration with energy density $\langle G^{2}\rangle\sim\Lambda_{QCD}^{4}$ spread over a “pancake” surface of thickness $\Lambda_{QCD}^{-2}$. This estimate interprets the dressing factor $\exp\left(-\kappa S[C]\right)$ as a statistical probability $\exp\left(-\beta\delta E\right)$ for a large quark loop propagating through the QCD vacuum (the dual Meissner effect).
@@ 13 Conclusion
This Hodge–dual minimal area $S[C]$ enters the non-perturbative factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop. This factor is an exact zero mode factor satisfying the MM equation in QCD [MMEq79, MM1981NPB, Mig83]. The QCD OPE provides the estimate (12.8) relating our mass scale $\kappa$ to the gluon condensate and QCD mass scale. The primary physical implication is a first-principles geometric proof of confinement, as stated in our Confinement Theorem (Theorem 9.1). This theorem provides an explicit, shape-dependent formula for the effective string tension $\sigma_{\mathrm{eff}}[C]$ (defined as a ratio of $S[C]$ to the Euclidean area $A[C]$). This parameter depends on the loop shape and is bounded from below as $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$. This solution lays the groundwork for a quantitative theory of the high-level hadron spectrum.